Please help me with the following question. Ireally need the help my work is due today and I just don't get it.
1.Set K contains 30 elements, set J contains 46 elements, and 11 elements are common to both sets. Find n ( K U J).
2.Given A = {1, 2, 3}, B = {3, 4, 5, 6,}, and C = {3, 5, 7}. Evaluate each set
a) A ∩ B
b) A ∩ C
c) A U C
d) B U C
e) (A U B) ∩ C
f) A U (B U C)
g) (A ∩ B) ∩ C
h) (A ∩ B) U C
Please help me

Answer by jim_thompson5910(21685)

# 1

To find the number of elements in the set K U J, simply add the two sets together. However, remember to subtract the number of elements in both sets (since these elements are repeated)


In other words, use this formula:

n ( K U J) = n(K) + n(J) - n(K ∩ J)


In this case, n(K)=30, n(J)=46 and n(K ∩ J) = 11. Plug these values in to get:

n ( K U J) = 30 + 46 - 11


Add

n ( K U J) = 76 - 11


Subtract
n ( K U J) = 65


So there are 65 elements in set K U J


-----------------------------------------------------

# 2

I'll do the first four to get you started

a)

A ∩ B = Set of elements that BOTH sets A and B have in common

A ∩ B = {1, 2, 3} ∩ {3, 4, 5, 6}

A ∩ B = {3}

----------------

b)

A ∩ C = Set of elements that BOTH sets A and C have in common

A ∩ C = {1, 2, 3} ∩ {3, 5, 7}

A ∩ C = {3}

----------------
c)

A U C = Combination of the sets A and C (remove any duplicates)

A U C = {1, 2, 3} U {3, 5, 7}

A U C = {1,2,3,5,7}

----------------
d)

B U C = Combination of the sets B and C (remove any duplicates)

B U C = {3, 4, 5, 6} U {3, 5, 7}

B U C = {3,4,5,6,7}